OpenStudy (anonymous):

Solve for x. 3 • 4 – 3 • 7 = 3(x – 7) A. 4 B. –3 C. 6 D. –10

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

What did you get?

OpenStudy (anonymous):

B

OpenStudy (igreen):

No, that's wrong. \(3 \cdot 4 - 3 \cdot 7 = 3(x - 7)\) \(12 - 21 = 3x - 21\) \(-9 = 3x - 21\) \(12= 3x \) Can you divide 12 / 3?

OpenStudy (anonymous):

4

OpenStudy (igreen):

Yep, so that's your answer.

OpenStudy (anonymous):

so A?